# Two classes of posets with real-rooted chain polynomials

### v2

@@ 1. Introduction
where $f_{i-1}(\Delta(P))$ is the number of $i$-element chains in $P$ (which are precisely the $(i-1)$-dimensional faces of $\Delta(P)$) and $n$ is the largest size of such a chain (which in one more than the dimension of $\Delta(P)$). For some purposes, one may focus on the corresponding $h$-polynomial
@@
where ${\mathfrak{S}}_{n}$ is the symmetric group of permutations of the set $[n]$ and for $w\in{\mathfrak{S}}_{n}$, ${\rm Des}(w)=\{i\in[n-1]:w(i)>w(i+1)\}$ and ${\rm des}(w)$ is the set and the number of descents of $w$, respectively. Thus, $A^{T}_{n}(x)$ is the descent enumerator for permutations in ${\mathfrak{S}}_{n}$ which have ascents at specified positions (those in $[n-1]{\smallsetminus}T$); it reduces to the classical Eulerian polynomial $A_{n}(x)$ [37, Section 1.4] for $T=[n-1]$. A $q$-analogue of $A^{T}_{n}(x)$ was studied in [17] in the special case $T=[r]\subseteq[n-1]$ (see also Example 3.4). To the best of our knowledge, except for the important special case of Eulerian polynomials, the unimodality, log-concavity and real-rootedness of $A^{T}_{n}(x)$ have not been considered before.
@@ 2.1. Polynomials
Given a polynomial $p(x)\in{\mathbb{R}}[x]$ of degree at most $n$, there exist unique symmetric polynomials $a(x),b(x)\in{\mathbb{R}}[x]$ with centers of symmetry $n/2$ and $(n-1)/2$, respectively, such that $p(x)=a(x)+xb(x)$. This expression is known as the symmetric decomposition (or Stapledon decomposition) of $p(x)$ with respect to $n$. Then, $p(x)$ is said to have a nonnegative (respectively, unimodal or real-rooted) symmetric decomposition with respect to $n$ if $a(x)$ and $b(x)$ have nonnegative coefficients (respectively, are unimodal or real-rooted); see [7, 14] for more information about these concepts. Every polynomial which has a nonnegative unimodal symmetric decomposition with respect to $n$ is unimodal, with a peak at position $\lceil n/2\rceil$.
@@ 2.2. Poset combinatorics
The zeta polynomial ${\mathcal{Z}}(P,x)$ is another important enumerative invariant of a finite poset $P$ [37, Section 3.12]. For the ease of notation, we define it here by letting ${\mathcal{Z}}(P,k)$ be the number of multichains $p_{1}\leq p_{2}\leq\cdots\leq p_{k}$ of length $k-1$ (rather than $k-2$) of elements of $P$, where ${\mathcal{Z}}(P,0):=1$. A comparison of [37, Proposition 3.12.1 (a)] with [34, Theorem II.1.4] then shows that
@@ 3. Permutations with restricted descent set
We proceed by induction on $n$, the result being trivial for $n=0$. Suppose that $n\geq 1$ and that (11) is an interlacing sequence of real-rooted polynomials when $n$ is replaced by $n-1$. It is straightforward to verify from the defining equation (9) that
@@
(a) For $T=[r]\subseteq[n-1]$, the polynomial $A^{T}_{n}(x)$ is the descent enumerator for permutations $w\in{\mathfrak{S}}_{n}$ which have ascents in the last $n-r-1$ positions. A $q$-analogue of $A^{T}_{n}(x)$ in this case was studied in [17] (although the unimodality of $A^{T}_{n}(x)$ was not addressed there). Theorem 3.1 and Corollary 3.2 imply that $A^{T}_{n}(x)$ is a real-rooted, hence unimodal, polynomial of degree $r$ and that it has a mode $m$ such that $\lfloor r/2\rfloor\leq m\leq\lceil(r+1)/2\rceil$. Since $(B_{n})_{T}$, with its maximum element removed, is a simplicial poset in this case, the real-rootedness of $A^{T}_{n}(x)$ already follows from the main result of [15]. Setting $q=1$ in the formula of [17, Theorem 2.10] gives that
@@
(b) More generally, for $T=\{s+1,s+2,\dots,s+r\}\subseteq[n-1]$, the polynomial $A^{T}_{n}(x)$ is the descent enumerator for permutations $w\in{\mathfrak{S}}_{n}$ which have ascents in the first $s$ and the last $n-r-s-1$ positions. According to Theorem 3.1 and Corollary 3.2, $A^{T}_{n}(x)$ is a real-rooted, hence unimodal, polynomial of degree $r$ which has a mode $m$ such that $\lfloor(r-1)/2\rfloor\leq m\leq\lceil(r+1)/2\rceil$. ∎
@@
Let $T=\{2,4,\dots,2n-2\}$, so that $A^{T}_{2n}(x)$ is the descent enumerator for permutations $w\in{\mathfrak{S}}_{2n}$ which have an ascent in every odd position. By Theorem 3.1 and Corollary 3.2, $A^{T}_{2n}(x)$ is a real-rooted, hence unimodal, polynomial of degree $n-1$ which has a mode $m_{n}$ such that $\lfloor 5(n-1)/6\rfloor\leq m_{n}\leq\lceil 5(n-1)/6\rceil$.
@@
We conclude this section with the following question. Part (b) provided a lot of the motivation behind this paper; it is an open problem [7, Question 7.2] to decide whether the inequalities which appear there hold for the $h$-vectors of all $(r-1)$-dimensional doubly Cohen–Macaulay simplicial complexes. An affirmative answer to part (a) would imply the (weaker) top-heavy inequalities $h_{i}(T)\leq h_{r-i}(T)$ for $0\leq i\leq\lfloor r/2\rfloor$; we refer the reader to [41] for this implication and for the concept of a convex ear decomposition.
@@ 4. Rank-selected subposets of simplicial posets
This section proves part (b) of Theorem 1.2 and gives an application. We recall that a finite poset $P$ with a minimum element $\hat{0}$ is said to be simplicial [33] [34, Section III.6] if the interval $[\hat{0},y]$ is isomorphic to a Boolean lattice for every $y\in P$. The enumerative invariant of a graded simplicial poset $P$ of rank $n$ which will be essential to the proof is the $h$-polynomial of $P$. This was defined by Stanley [33] as
@@
for $j\in\{0,1,\dots,n\}$ and $h(P,x)=h(\Delta,x)$, if $P$ is the face poset of an $(n-1)$-dimensional simplicial complex $\Delta$. Stanley [33] (see also [34, Section III.6]) showed that $h(P,x)$ has nonnegative coefficients for every Cohen–Macaulay simplicial poset $P$.
@@
for every $T\subseteq[n]$. Let us write $T=\{a_{1},a_{2},\dots,a_{r}\}\subseteq[n]$, with $1\leq a_{1}<\cdots<a_{r}\leq n$. There are $f_{a_{r}-1}(P)$ elements of rank $a_{r}$ in $P$ and (${a_{r}\choose a_{1},a_{2}-a_{1},\dots,a_{r}-a_{r-1}}$ chains of elements of ranks $a_{1},a_{2},\dots,a_{r-1}$ in any Boolean lattice of rank $a_{r}$. Given this and Equation (14), we find that
@@
Thus, it suffices to verify that (${n-k\choose a_{r}-k}{a_{r}\choose a_{1},a_{2}-a_{1},\dots,a_{r}-a_{r-1}}$ is equal to the number of permutations $w\in{\mathfrak{S}}_{n+1}$ such that $w(n+1)=k+1$ and ${\rm Asc}(w)\subseteq T$, a task which can safely be left to the reader. ∎
@@
We will apply Theorem 1.2 to the poset of $r$-colored subsets of the set $[n]$, defined as follows. We consider the subsets $\Omega$ of $[n]\times\{0,1,\dots,r-1\}$ for which for every $i\in[n]$ there is at most one $j\in\{0,1,\dots,r-1\}$ such that $(i,j)\in\Omega$ and let $P$ be the set of all such subsets, partially ordered by inclusion. Thus, $P$ is a graded simplicial poset of rank $n$ which is isomorphic to the Boolean lattice $B_{n}$ for $r=1$. It was shown in the proof of [2, Theorem 1.3] that $P$ is shellable, hence Cohen–Macaulay, and that $\beta_{\hat{P}}(S)$ is equal to the number of $r$-colored permutations $u\in{\mathfrak{S}}_{n}[{\mathbb{Z}}_{r}]$ with descent set equal to $S$, for every $S\subseteq[n]$. As a result, in view of Equation (7),
@@ 5. Noncrossing partition lattices
This section proves Theorem 1.3. We first recall the definition of ${\rm NC}_{W}$. Let $W$ be a finite Coxeter group with rank $r_{W}$ and set of reflections ${\mathcal{T}}$. For $\alpha\in W$ we denote by $\ell_{\mathcal{T}}(\alpha)$ the smallest $k$ such that $\alpha$ can be written as a product of $k$ reflections in ${\mathcal{T}}$. We define the partial order $\preceq$ on $W$ by letting $\alpha\preceq\beta$ if $\ell_{\mathcal{T}}(\alpha)+\ell_{\mathcal{T}}(\alpha^{-1}\beta)=\ell_{\mathcal{T}}(\beta)$, in other words if there exists a shortest factorization of $\alpha$ into reflections which is a prefix of such a shortest factorization of $\beta$. Then, ${\rm NC}_{W}$ is defined as the closed interval $[e,\gamma]$ in $(W,\preceq)$, where $e\in W$ is the identity element and $\gamma$ is any Coxeter element of $W$. The noncrossing partition poset ${\rm NC}_{W}$ is a rank-symmetric, graded lattice with rank function $\ell_{\mathcal{T}}$ and rank $r_{W}$; its combinatorial type is independent of the choice of $\gamma$. A detailed exposition of noncrossing partition lattices can be found in [1, Chapter 2].
@@
$h(\Delta({\rm NC}_{W}),x)=\begin{cases}1+(m-1)x,&\text{if ${\mathcal{X}}=I_{2}(m)$}\\ 1+28x+21x^{2},&\text{if ${\mathcal{X}}=H_{3}$}\\ 1+275x+842x^{2}+232x^{3},&\text{if ${\mathcal{X}}=H_{4}$}\\ 1+100x+265x^{2}+66x^{3},&\text{if ${\mathcal{X}}=F_{4}$}\\ 1+826x+10778x^{2}+21308x^{3}+8141x^{4}+418x^{5},&\text{if ${\mathcal{X}}=E_{6}$}\\ 1+4152x+110958x^{2}+446776x^{3}+412764x^{4}&\\ \ \ +\ 85800x^{5}+2431x^{6},&\text{if ${\mathcal{X}}=E_{7}$}\\ 1+25071x+1295238x^{2}+9523785x^{3}+17304775x^{4}&\\ \ \ +8733249x^{5}+1069289x^{6}+17342x^{7},&\text{if ${\mathcal{X}}=E_{8}$}.\end{cases}$
@@
where the first two sums run through all compositions $(a_{1},a_{2}\dots,a_{k+1})$ of $n$ with $k+1$ parts. Using the fact that (${n-2\choose a_{i}-2}=\frac{a_{i}-1}{n-1}{n-1\choose a_{i}-1}$, changing the order of summation in the double sum and replacing $a_{i}$ with $a_{i}+1$ yields that
@@
where ${\mathcal{W}}_{n,r}$ is the set of words $w:\{0,1,\dots,n\}\rightarrow[r]$ with $w(0)=1$ and ${\rm asc}^{*}(w)$ is the number of indices $i\in[n]$ such that $w(i-1)<w(i)$. We note that
$x^{n-1}E_{n,r}(1/x)=\sum_{w\in[r]^{n}}x^{n-1-{\rm des}(w)}=\sum_{w\in[r]^{n}}x^{{\rm asc}^{*}(w)},$
where ${\rm asc}^{*}(w)=n-1-{\rm des}(w)$ is the number of strict ascents of $w\in[r]^{n}$. Distinguishing the cases $w(1)=1$ and $w(1)\geq 2$ for such a word and for a word $w\in{\mathcal{W}}_{n,r}$ we get
$x^{n-1}E_{n,r}(1/x)$ $=\tilde{E}_{n-1,r}(x)+\sum_{w\in[r]^{n}:\,w(1)\geq 2}x^{{\rm asc}^{*}(w)},$
$\tilde{E}_{n,r}(x)$ $=\tilde{E}_{n-1,r}(x)+x\sum_{w\in[r]^{n}:\,w(1)\geq 2}x^{{\rm asc}^{*}(w)}.$

### v3

@@ 1. Introduction
where $f_{i-1}(\Delta(P))$ is the number of $i$-element chains in $P$ (which are precisely the (i1)(i-1)-dimensional faces of $\Delta(P)$) and $n$ is the largest size of such a chain (which is one more than the dimension of $\Delta(P)$). For some purposes, one may focus on the corresponding $h$-polynomial
@@
where ${\mathfrak{S}}_{n}$ is the symmetric group of permutations of the set $[n]$ and for $w\in{\mathfrak{S}}_{n}$, ${\rm Des}(w)=\{i\in[n-1]:w(i)>w(i+1)\}$ and ${\rm des}(w)$ is the set and the number of descents of $w$, respectively. Thus, $A^{T}_{n}(x)$ is the descent enumerator for permutations in ${\mathfrak{S}}_{n}$ which have ascents at specified positions (those in $n-1]{\smallsetminus}T$); it reduces to the classical Eulerian polynomial $A_{n}(x)$ [37, Section 1.4] for T=[n1]T=[n-1]. A $q$-analogue of $A^{T}_{n}(x)$ was studied in [17] in the special case $T=[r]\subseteq[n-1]$ (see also Example 3.4). To the best of our knowledge, except for the important special case of Eulerian polynomials, the unimodality, log-concavity and real-rootedness of $A^{T}_{n}(x)$ have not been considered before.
@@ 2.1. Polynomials
Given a polynomial $p(x)\in{\mathbb{R}}[x]$ of degree at most $n$, there exist unique symmetric polynomials $a(x),b(x)\in{\mathbb{R}}[x]$ with centers of symmetry $n/2$ and (n1)/2(n-1)/2, respectively, such that $p(x)=a(x)+xb(x)$. This expression is known as the symmetric decomposition (or Stapledon decomposition) of $p(x)$ with respect to $n$. Then, $p(x)$ is said to have a nonnegative (respectively, unimodal or real-rooted) symmetric decomposition with respect to $n$ if $a(x)$ and $b(x)$ have nonnegative coefficients (respectively, are unimodal or real-rooted); see [7, 14] for more information about these concepts. Every polynomial which has a nonnegative unimodal symmetric decomposition with respect to $n$ is unimodal, with a peak at position $\lceil n/2\rceil$.
@@ 2.2. Poset combinatorics
The zeta polynomial ${\mathcal{Z}}(P,x)$ is another important enumerative invariant of a finite poset $P$ [37, Section 3.12]. For the ease of notation, we define it here by letting ${\mathcal{Z}}(P,k)$ be the number of multichains $p_{1}\leq p_{2}\leq\cdots\leq p_{k}$ of length k1k-1 (rather than k2k-2) of elements of $P$, where ${\mathcal{Z}}(P,0):=1$. A comparison of [37, Proposition 3.12.1 (a)] with [34, Theorem II.1.4] then shows that
@@ 3. Permutations with restricted descent set
We proceed by induction on $n$, the result being trivial for $n=0$. Suppose that $n\geq 1$ and that (11) is an interlacing sequence of real-rooted polynomials when $n$ is replaced by n1n-1. It is straightforward to verify from the defining equation (9) that
@@
(a) For $T=[r]\subseteq[n-1]$, the polynomial $A^{T}_{n}(x)$ is the descent enumerator for permutations $w\in{\mathfrak{S}}_{n}$ which have ascents in the last nr1n-r-1 positions. A $q$-analogue of $A^{T}_{n}(x)$ in this case was studied in [17] (although the unimodality of $A^{T}_{n}(x)$ was not addressed there). Theorem 3.1 and Corollary 3.2 imply that $A^{T}_{n}(x)$ is a real-rooted, hence unimodal, polynomial of degree $r$ and that it has a mode $m$ such that $\lfloor r/2\rfloor\leq m\leq\lceil(r+1)/2\rceil$. Since $(B_{n})_{T}$, with its maximum element removed, is a simplicial poset in this case, the real-rootedness of $A^{T}_{n}(x)$ already follows from the main result of [15]. Setting $q=1$ in the formula of [17, Theorem 2.10] gives that
@@
(b) More generally, for $T=\{s+1,s+2,\dots,s+r\}\subseteq[n-1]$, the polynomial $A^{T}_{n}(x)$ is the descent enumerator for permutations $w\in{\mathfrak{S}}_{n}$ which have ascents in the first $s$ and the last nrs1n-r-s-1 positions. According to Theorem 3.1 and Corollary 3.2, $A^{T}_{n}(x)$ is a real-rooted, hence unimodal, polynomial of degree $r$ which has a mode $m$ such that $\lfloor(r-1)/2\rfloor\leq m\leq\lceil(r+1)/2\rceil$. ∎
@@
Let $T=\{2,4,\dots,2n-2\}$, so that $A^{T}_{2n}(x)$ is the descent enumerator for permutations $w\in{\mathfrak{S}}_{2n}$ which have an ascent in every odd position. By Theorem 3.1 and Corollary 3.2, $A^{T}_{2n}(x)$ is a real-rooted, hence unimodal, polynomial of degree n1n-1 which has a mode $m_{n}$ such that $\lfloor 5(n-1)/6\rfloor\leq m_{n}\leq\lceil 5(n-1)/6\rceil$.
@@
We conclude this section with the following question. Part (b) provided a lot of the motivation behind this paper; it is an open problem [7, Question 7.2] to decide whether the inequalities which appear there hold for the $h$-vectors of all (r1)(r-1)-dimensional doubly Cohen–Macaulay simplicial complexes. An affirmative answer to part (a) would imply the (weaker) top-heavy inequalities $h_{i}(T)\leq h_{r-i}(T)$ for $0\leq i\leq\lfloor r/2\rfloor$; we refer the reader to [41] for this implication and for the concept of a convex ear decomposition.
@@ 4. Rank-selected subposets of simplicial posets
This section proves part (b) of Theorem 1.2 and gives an application. We recall that a finite poset $P$ with a minimum element $\hat{0}$ is said to be simplicial [33] [34, Section III.6] if the interval $\hat{0},y]$ is isomorphic to a Boolean lattice for every $y\in P$. The enumerative invariant of a graded simplicial poset $P$ of rank $n$ which will be essential to the proof is the $h$-polynomial of $P$. This was defined by Stanley [33] as
@@
for $j\in\{0,1,\dots,n\}$ and $h(P,x)=h(\Delta,x)$, if $P$ is the face poset of an (n1)(n-1)-dimensional simplicial complex $\Delta$. Stanley [33] (see also [34, Section III.6]) showed that $h(P,x)$ has nonnegative coefficients for every Cohen–Macaulay simplicial poset $P$.
@@
for every $T\subseteq[n]$. Let us write $T=\{a_{1},a_{2},\dots,a_{r}\}\subseteq[n]$, with $1\leq a_{1}<\cdots<a_{r}\leq n$. There are $f_{a_{r}-1}(P)$ elements of rank $a_{r}$ in $P$ and $a_{r}\choose a_{1},a_{2}-a_{1},\dots,a_{r}-a_{r-1}$} chains of elements of ranks $a_{1},a_{2},\dots,a_{r-1}$ in any Boolean lattice of rank $a_{r}$. Given this and Equation (14), we find that
@@
Thus, it suffices to verify that ${n-k\choose a_{r}-k}{a_{r}\choose a_{1},a_{2}-a_{1},\dots,a_{r}-a_{r-1}}$ is equal to the number of permutations $w\in{\mathfrak{S}}_{n+1}$ such that $w(n+1)=k+1$ and ${\rm Asc}(w)\subseteq T$, a task which can safely be left to the reader. ∎
@@
We will apply Theorem 1.2 to the poset of $r$-colored subsets of the set $[n]$, defined as follows. We consider the subsets $\Omega$ of $n]\times\{0,1,\dots,r-1\}$ for which for every $i\in[n]$ there is at most one $j\in\{0,1,\dots,r-1\}$ such that $(i,j)\in\Omega$ and let $P$ be the set of all such subsets, partially ordered by inclusion. Thus, $P$ is a graded simplicial poset of rank $n$ which is isomorphic to the Boolean lattice $B_{n}$ for $r=1$. It was shown in the proof of [2, Theorem 1.3] that $P$ is shellable, hence Cohen–Macaulay, and that $\beta_{\hat{P}}(S)$ is equal to the number of $r$-colored permutations $u\in{\mathfrak{S}}_{n}[{\mathbb{Z}}_{r}]$ with descent set equal to $S$, for every $S\subseteq[n]$. As a result, in view of Equation (7),
@@ 5. Noncrossing partition lattices
This section proves Theorem 1.3. We first recall the definition of ${\rm NC}_{W}$. Let $W$ be a finite Coxeter group with rank $r_{W}$ and set of reflections $\mathcal{T}$}. For $\alpha\in W$ we denote by $\ell_{\mathcal{T}}(\alpha)$ the smallest $k$ such that $\alpha$ can be written as a product of $k$ reflections in $\mathcal{T}$}. We define the partial order $\preceq$ on $W$ by letting $\alpha\preceq\beta$ if $\ell_{\mathcal{T}}(\alpha)+\ell_{\mathcal{T}}(\alpha^{-1}\beta)=\ell_{\mathcal{T}}(\beta)$, in other words if there exists a shortest factorization of $\alpha$ into reflections which is a prefix of such a shortest factorization of $\beta$. Then, ${\rm NC}_{W}$ is defined as the closed interval $e,\gamma]$ in $(W,\preceq)$, where $e\in W$ is the identity element and $\gamma$ is any Coxeter element of $W$. The noncrossing partition poset ${\rm NC}_{W}$ is a rank-symmetric, graded lattice with rank function $\ell_{\mathcal{T}}$ and rank $r_{W}$; its combinatorial type is independent of the choice of $\gamma$. A detailed exposition of noncrossing partition lattices can be found in [1, Chapter 2].
@@
[…]
@@
where the first two sums run through all compositions $(a_{1},a_{2}\dots,a_{k+1})$ of $n$ with $k+1$ parts. Using the fact that ${n-2\choose a_{i}-2}=\frac{a_{i}-1}{n-1}{n-1\choose a_{i}-1}$, changing the order of summation in the double sum and replacing $a_{i}$ with $a_{i}+1$ yields that
@@
where ${\mathcal{W}}_{n,r}$ is the set of words $w:\{0,1,\dots,n\}\rightarrow[r]$ with $w(0)=1$ and ${\rm asc}^{*}(w)$ is the number of indices $i\in[n]$ such that w(i1)<w(i)w(i-1)<w(i). We note that
$x^{n-1}E_{n,r}(1/x)=\sum_{w\in[r]^{n}}x^{n-1-{\rm des}(w)}=\sum_{w\in[r]^{n}}x^{{\rm asc}(w)},$
where ${\rm asc}(w)=n-1-{\rm des}(w)$ is the number of strict ascents of $w\in[r]^{n}$. Distinguishing the cases $w(1)=1$ and $w(1)\geq 2$ for such a word and for a word $w\in{\mathcal{W}}_{n,r}$ we get
$x^{n-1}E_{n,r}(1/x)$ $=\tilde{E}_{n-1,r}(x)+\sum_{w\in[r]^{n}:\,w(1)\geq 2}x^{{\rm asc}(w)},$
$\tilde{E}_{n,r}(x)$ $=\tilde{E}_{n-1,r}(x)+x\sum_{w\in[r]^{n}:\,w(1)\geq 2}x^{{\rm asc}(w)}.$